# Schelling segregation dynamics in densely-connected social network graphs

### v1

@@ Abstract
We found that a dense social network does not become as strongly segregated as a sparse network, and that agents’ numbers of same-group neighbours do not greatly exceed their desired numbers (i.e. they do not end up more segregated than they desire to be). Furthermore, we found that the network was very difficult to polarise when one group was somewhat smaller than the other, and that it became unstable when one group was extremely small, both of which provide insights into real-world polarisation dynamics. Finally, we tested the question of whether an increase in the minority group’s desire for same-group neighbours created more segregation than a similar increase for the majority group — the ”paradox of weak minority preferences” — and found mixed evidence for this effect in a densely connected social network.
@@ Introduction
We are interested in developing methods of measuring polarisation that do not focus on political parties. This is because we believe that the polarising division tends to occur around issues that do not map directly to political party stances, and may not be captured well by voter data. For example, voter data in the USA does not capture the views of Black people before 1964, and analyses which use it are therefore missing an important part of the society (Hetherington, (2009)). Furthermore, methods which focus on political parties require sorting the parties into a left-right ideological divide, and this is difficult in the New Zealand case. The populist New Zealand First party frequently receives enough of the vote to be placed in a ”kingmaker” position in coalition negotiations, and has entered coalitions with both left- and right-wing parties during its history; in 2017 it formed a coalition with the left-wing Labour party, and in 2023 it formed a coalition with the right-wing National and Act parties. This means that changes in New Zealand’s political environment cannot be captured just by a left-right divide, and the political field cannot be simplified by excluding New Zealand First due to their frequent importance in coalition formation. Furthermore, there are several unique factors in the New Zealand political landscape that affect the country’s susceptibility to populist politics (Vowles and Curtin, (2020)), which means that we may benefit from new methods to detect populism.
@@
Our previous work demonstrated that measuring the embedded dimensionality $\hat{d}$ of a social network, as represented by a Random Dot Product Graph, may be a useful indicator of whether that network is polarised (Anastasi and Dalla Riva, 2025a ). This is because we define a ”polarised” social network as one which is split into two groups with very few or no connections between them; in real life, examples of this are the policies of South African apartheid or US-American racial segregation, in which people belonging to the two groups are prevented by law and violence from making social connections with each other. When expressed as a Random Dot Product Graph, a polarised network will have a low $\hat{d}$, and $\hat{d}$ will increase as connections are made between the groups. We do not suggest that there is a specific value of the embedded dimensionality at which a network becomes polarised, but rather that observation of a network over time can show polarisation occurring if the embedded dimensionality decreases over time. This can also work as a method for detecting populism if it is based on creating a deep divide between the populist group and an out-group; the political science work we draw from refers to polarisation as ”populist antagonism” (Laclau and Mouffe, (1985)).
@@
We have turned to Schelling segregation as a straightforward way of investigating some of the dynamics underlying polarisation. If a polarised society is one where members of the two groups are prevented from making connections, then the dynamics of agents wanting certain numbers of same-group connections are likely to be meaningful for whether a network is polarised or not. The Schelling segregation model was established in 1971, with simple rules dictating agents’ behaviour on a grid (Schelling, (1971)). Agents were ”happy” when a sufficient number of their neighbours were in the same group; if not, they moved to a vacant space. Schelling found that even low levels of same-neighbour preferences resulted in patterns of segregation. Using a Schelling segregation model in a social network gives us the opportunity to further test some assumptions about polarisation which we have already explored using stochastic block models (Anastasi and Dalla Riva, 2025a ). The first assumption is that polarisation decreases when the groups are more connected, and the second is that polarisation decreases as one group becomes larger than the other. We found clear support for the first, and mixed support for the second; while polarisation decreased as one group became large, when it was much larger than the other (99-to-1) polarisation rapidly increased again. In a Schelling model, this first assumption corresponds to the agents’ minimum number of same-group neighbours needed to be happy, and the second one corresponds to the sizes of the groups.
@@
Many extensions to the original model have been made.Clark, (1991) evaluated the Schelling model using real same-group preference data from surveys in the USA, and found that while the real preferences did not match Schelling’s assumptions they still generated segregation. Gilbert, (2002) uses variations on the Schelling model to show that higher-order effects which match residential segregation can be found from many types of models, so models of residential segregation must be validated at both the individual and macro level. Zhang, (2004) found that when the initial model is segregated rather than random, white agents will leave neighbourhoods as black agents enter it, preserving the segregation even when it is not optimal for their overall preferences regarding amenities etc. Fossett and Waren, (2005) found that while a 50-50 ethnic mix is often used in simulation, other ethic mixes have notable effects on the results and a 50-50 mix may not apply to many real-world scenarios. They also find that a small added preference for neighbours of the other group reduces the segregation effect compared to neutrality towards the other group. Fossett, (2006) raises the issue that if the minority group want to be ”more integrated” than their presence in the city allows (e.g. wanting a 50-50 neighbourhood despite an overall 80-20 ethnic mix), this can produce segregation by pulling other members of the minority into one specific neighbourhood to satisfy this preference, meaning that the ethnic mix is reduced everywhere else; as such, existing patterns of segregation can be maintained even by the tactics used by the minority to try and protect themselves. Clark and Fossett, (2008) continue this line of inquiry and find that the unwillingness of any group to be a minority in ”their own neighbourhood” can maintain segregation even in the absence of animosity towards other groups, even when additional dynamics such as housing cost are considered. Fagiolo et al., (2007) show that segregation occurs in several types of spaces, such as non-directed graphs, indicating that Schelling segregation is not due to the topological constraints of the grid space used in the original model. Similarly, Domic et al., (2011) used lattices in two and three dimensions to explore further spatial elements of the model and whether thermodynamic principles of energy can be used to model the emergent segregation dynamics. Radi and Gardini, (2015) found that limiting how many of the minority group can move into a given neighbourhood at once helped promote integration, by preventing the influx from causing major destabilisation. Gandica et al., (2016) found that the topology of the space affects the early stages of the model at high tolerance values, but at lower tolerance values the segregation process is topology-independent. Collard and Ghetiu, (2016) added a parameter to measure how segregated each individual was from the other group, so that more segregated agents were higher vertically in the resulting plots; examining their plots as if they were topographical maps, they found that models with low other-group tolerance were segregated into ”hills and valleys” rather than the more even ”ruggedness” of higher-tolerance models. Jani, (2020) tested the effects of housing scarcity on Schelling segregation, finding that segregation initially increased as resources became more scarce, before finally collapsing when resources were too scarce to sustain it. Ishida, (2024) explored the role of fuzzy group membership (i.e. membership as a continuous variable rather than a discrete one) on segregation patterns; they found that less-fuzzy agents tended to be more unhappy and ”wander” for longer, and when they finally settled it was often near agents of the other group (since spaces near more similar agents had already been occupied), which they suggest offers another potential mechanism for explaining ethnic conflict in a neighbourhood. Blair, (2023) found that whether white agents departed from a neighbourhood was related to the cost of outside options, including that higher desire to be near other white agents could prevent an agent from leaving the neighbourhood as it integrates, rather than departing as is commonly expected.
@@
We are interested in expanding Schelling segregation to a network model where the agents are connected directly, rather than using a network graph as the space in which agents move. Modern social networks are less dependent on geographic proximity than they have been in the past, so a person in an integrated neighbourhood may still be able to live a highly segregated life, provided that their other social connections are sufficiently segregated. As such, we give particular attention to the following two papers which use a network graph in this way. Henry et al., (2011) use a Markov chain to generate a stochastic sequence of graphs comprised of actors (nodes) connected by links (edges). The model evolves based on Schelling-based assessments of the attribute distance between two actors connected by a given edge, with links being terminated and rewired to prioritise connections between actors with smaller attribute distances. They use networks of 100 actors connected by 300 edges, and find that segregation emerges in the graps in similar ways to spatial Schelling models. Similarly, Gretha et al., (2018) create graphs of 1000 agents with 4 initial links, and base network evolution behaviours on a tolerance of 50%. They find that segregation emerges in the network that is greater than the amount desired by the agents, similar to in Schelling’s original results (though this may not be surprising given that this effect has been observed with tolerances as low as 37.5%). Curiously, random behaviour in the selection of which agent to make a new link with caused cascades of dissatisfaction and more movement within the network, whereas ”discriminatory” selection of an agent in the same group caused fewer cascades – an important result for understanding how the dynamics of networks are different to those of spatial models. Based on these results we believe that there is scope for new research that uses much larger and more densely connected graphs, in order to model an agent’s social life more extensively. One of our aims with this paper is to address this gap.
@@
Secondly, we systematically vary both the tolerance of the two groups and their proportions. The combination of these variables has been found to be very important to the Schelling model. For example, Fossett, (2006) has found that if the small group wish to have a higher proportion of same-group neighbours than exist in the overall population, the model will tend towards segregation as minority group members congregate to achieve happiness (since above-average numbers of a minority in one area naturally lead to below-average numbers in another). While a social network is not subject to the spatial considerations that have previously applied to these results, it is still possible that this interaction will affect a social network model, as minority group members may become inundated with links to majority group members, making it difficult to achieve the desired ratio of same- and different-group connections. However, this must be tested experimentally rather than presumed.
Finally, we measure $\hat{d}$ alongside the ratios of same- and different-group connections for each agent, as a further investigation of whether it is a good measurement of polarisation in social networks. Our preliminary results indicate that $\hat{d}$ increases when there are more connections between two groups – i.e. the groups are less polarised – and that it decreases sharply when one group is much larger than the other. Since existing research on the Schelling model has found that segregation and agent unhappiness are both high when one group is much larger than the other, we are interested to see what results we get by measuring $\hat{d}$ for the networks generated in our experiment.
@@ 1.1 Agent-based model
We designed an agent-based model in Julia that uses the preference rules of Schelling segregation to change the networking behaviour of the agents. It takes parameters $t\in[0,1]$ for tolerance ($t_{1}\in[0,1]$ and $t_{2}\in[0,1]$ when the tolerance is different between the groups), i.e. the proportion of same-group neighbours needed for the agent to be happy, and $s\in[0,1]$ for the size of the smaller group. The model is initialised with 1000 agents split between 2 groups, who each make 80 random connections on a simple weighted network graph. At each step, each agent evaluates how many of its network neighbours are in the same group as it. If the proportion of same-group neighbours is equal to or above its tolerance $t$, it is happy and does nothing. If the proportion of same-group neighbours is below $t$, it breaks a connection with a neighbour of the opposite group. If at any point the number of total neighbours is lower than 40, it adds a neighbour at random. The model runs for 1000 steps or until all agents are happy. All models were repeated 100 times for each combination of parameters.
@@ 1.2 Network modelling
In order to measure the embedded dimensionality $\hat{d}$ we model the simple weighted network graphs as Random Dot-Product Graphs (RDPGs). This graph embedding is chosen because its optimal embedded dimensionality is established before the analysis and is independent of the network size. For a full discussion of the use of RDPG, see Anastasi and Dalla Riva, 2025a .
@@ 1.3 Embedding dimension
We define the dimension of a communication network as the optimal choice of $d$ for the RDPG embedding of the network and denote it $\hat{d}$.
@@
An RDPG consiting of entirely random connections will have a $\hat{d}$ of 1, since 1 dimension is required for the nodes to make connections at all. It will not have a higher dimensionality because there is nothing else affecting how the edges are distributed. As such, instead of observing a high $\hat{d}$ which decreases as the network becomes segregated, as we have found in our previous work, the network will have a $\hat{d}$ of 1 until it is segregated enough that this noticeably affects whether edges are able to form between the two groups, at which point $\hat{d}$ will increase to 2. We note this because in our previous work we have used stochastic block models, which have enough structure to have a higher $\hat{d}$, whereas the random simple weighted graphs used in this paper are much more random and therefore start with a $\hat{d}$ of 1.
@@ 2.3 Changing group size
To investigate the effect of changing group size, we ran the same model with symmetrical tolerances of $t=0.25$, $t=0.5$, $t=0.85$, and $t=0.95$ so that group size effects were captured when the model was polarised and when it was not. For each tolerance, a minority group size $0.05\leq s\leq 0.5$ was tested in increments of 0.05. We could not test $s\leq 0.05$ due to the minimum of 40 neighbours set for the agent; since $s=0.01$ would give only 10 agents in the minority, they would be structurally incapable of being happy if they have $t_{1}\geq 0.25$ and the model would be unable to stabilise at all. We measured the number of steps it took for all agents to become happy (fig7), the proportion of similar neighbours to total neighbours (fig8), and $\hat{d}$ (fig9), with full results reported in Table 3.
@@
Figure 9: Boxplot of the embedded dimensionality of the networks at each group size.

### v2

@@ Abstract
We found that a dense social network does not become as strongly segregated as a sparse network, and that agents’ numbers of same-group neighbours do not greatly exceed their desired numbers (i.e. they do not end up more segregated than they desire to be). Furthermore, we found that the network was very difficult to polarise when one group was somewhat smaller than the other, and that the network became unstable when one group was extremely small; both phenomena may help explain the complexity of real-world polarisation dynamics, such as unique risks faced by very small groups in a society. Finally, we tested Fossett’s Fossett, (2006) “paradox of weak minority preferences”, a well-established result in grid- and map-based models which shows that an increase in the minority group’s desire for same-group neighbours can create more segregation than a similar increase for the majority group. In a densely connected social network, we find that the evidence for this effect is mixed.
@@ Introduction
We are interested in developing methods of measuring polarisation that do not focus on political parties. This is because we believe that the polarising division may occur around issues that do not map directly to political party stances, and may not be captured well by voter data. For example, voter data in the USA does not capture the views of Black people before 1964, and analyses which use it are therefore missing an important part of the society (Hetherington, (2009)). Furthermore, methods which focus on political parties require sorting the parties into a left-right ideological divide, and this is difficult in the New Zealand case. The populist New Zealand First party frequently receives enough of the vote to be placed in a ”kingmaker” position in coalition negotiations, and has entered coalitions with both left- and right-wing parties during its history; in 2017 it formed a coalition with the left-wing Labour party, and in 2023 it formed a coalition with the right-wing National and Act parties. This means that changes in New Zealand’s political environment cannot be captured just by a left-right divide, and the political field cannot be simplified by excluding New Zealand First due to their frequent importance in coalition formation. Furthermore, there are several unique factors in the New Zealand political landscape that affect the country’s susceptibility to populist politics (Vowles and Curtin, (2020)), which means that we may benefit from new methods to detect populism.
@@
Our previous work developed an indicator of polarisation based on the embedded dimensionality $\hat{d}$ of a network of connections between the users, i.e. the number of SVD dimensions required to capture the network interactions, when the network is represented by a Random Dot Product Graph (Anastasi and Dalla Riva, 2025a ). We refer to this network of connections as the ”social network”, meaning the network graph, rather than a ”social networking site” such as Facebook. We do not suggest that there is a specific value of the embedded dimensionality at which a network becomes polarised, but rather that observation of a network over time can show polarisation occurring if the embedded dimensionality decreases over time. This can also work as a method for detecting populism if it is based on creating a deep divide between the populist group and an out-group; the political science work we draw from refers to polarisation as ”populist antagonism” (Laclau and Mouffe, (1985)). When expressed as a Random Dot Product Graph, a polarised network will have a low $\hat{d}$, and $\hat{d}$ will increase as connections are made between the groups. In the abstract, a ”maximally polarised” social network would be one which is split into two groups with no connections between them at all, giving a $\hat{d}$ of 2. In real life this is not achievable due to the complexity of nation-sized social systems, but we believe that it is still possible for a nation to pursue a minimal $\hat{d}$ through systems of law and violence that are intended to create two separate poles. Examples of national systems we would consider polarised are South Africa under apartheid and the USA during the early 20th Century era of racial segregation.
@@
We have turned to Schelling segregation as a straightforward way of investigating some of the dynamics underlying polarisation. If a polarised society is one where members of the two groups are prevented from making connections, then the dynamics of agents wanting certain numbers of same-group connections are likely to be meaningful for whether a network is polarised or not. The Schelling segregation model was established in 1971, with simple rules dictating agents’ behaviour on a grid (Schelling, (1971)). Agents were ”happy” when a sufficient number of their neighbours were in the same group; if not, they moved to a vacant space. Schelling found that even low levels of same-neighbour preferences resulted in patterns of segregation. Using a Schelling segregation model in a social network gives us the opportunity to further test some hypotheses about polarisation which we have already explored using Stochastic Block Models (Anastasi and Dalla Riva, 2025a ). The first hypothesis is that polarisation decreases when the groups are more connected, and the second is that polarisation decreases as one group becomes larger than the other. We found clear support for the first, and mixed support for the second; while polarisation decreased as one group became large, when it was much larger than the other (99-to-1) polarisation rapidly increased again. In a Schelling model, the critical parameter at play in for the first hypothesis is the agents’ minimum number of same-group neighbours needed to be happy, and for the second hypothesis the critical parameter is the sizes of the groups.
@@
Many extensions to the original model have been made.Clark, (1991) evaluated the Schelling model using real same-group preference data from surveys in the USA, and found that while the real preferences did not match Schelling’s assumptions they still generated segregation. Gilbert, (2002) uses variations on the Schelling model to show that higher-order effects which match residential segregation can be found from many types of models, so models of residential segregation must be validated at both the individual and macro level. Zhang, (2004) found that when the initial model is segregated rather than random, white-ethnicity agents will leave neighbourhoods as black-ethnicity agents enter them, preserving the segregation even when it is not optimal for their overall preferences regarding amenities etc. Fossett and Waren, (2005) found that while a 50-50 ethnic mix is often used in simulation, other ethic mixes have notable effects on the results and a 50-50 mix may not apply to many real-world scenarios. They also find that a small added preference for neighbours of the other group reduces the segregation effect compared to neutrality towards the other group. Fossett, (2006) raises the ”paradox of weak minority preferences”, where the minority group can be the drivers of segregation if they want to be ”more integrated” than their presence in the city allows (e.g. wanting a 50-50 neighbourhood despite an overall 80-20 ethnic mix). This minority preference pulls other members of the minority into one specific neighbourhood to satisfy their happiness requirements, which reduces the ethnic mix everywhere else on the map; as such, existing patterns of segregation can be maintained even by the tactics used by the minority to try and protect themselves. Clark and Fossett, (2008) continue this line of inquiry and find that the unwillingness of any group to be a minority in ”their own neighbourhood” can maintain segregation even in the absence of animosity towards other groups, even when additional dynamics such as housing cost are considered. Fagiolo et al., (2007) show that segregation occurs in several types of spaces, such as non-directed graphs, indicating that Schelling segregation is not due to the topological constraints of the grid space used in the original model. Similarly, Domic et al., (2011) used lattices in two and three dimensions to explore further spatial elements of the model and whether thermodynamic principles of energy can be used to model the emergent segregation dynamics. Radi and Gardini, (2015) found that limiting how many of the minority group can move into a given neighbourhood at once helped promote integration, by preventing the influx from causing major destabilisation. Gandica et al., (2016) found that the topology of the space affects the early stages of the model at high tolerance values, but at lower tolerance values the segregation process is topology-independent. Collard and Ghetiu, (2016) added a parameter to measure how segregated each individual was from the other group, so that more segregated agents were higher vertically in the resulting plots; examining their plots as if they were topographical maps, they found that models with low other-group tolerance were segregated into ”hills and valleys” rather than the more even ”ruggedness” of higher-tolerance models. Jani, (2020) tested the effects of housing scarcity on Schelling segregation, finding that segregation initially increased as resources became more scarce, before finally collapsing when resources were too scarce to sustain it. Ishida, (2024) explored the role of fuzzy group membership (i.e. membership as a continuous variable rather than a discrete one) on segregation patterns; they found that less-fuzzy agents tended to be more unhappy and ”wander” for longer, and when they finally settled it was often near agents of the other group (since spaces near more similar agents had already been occupied), which they suggest offers another potential mechanism for explaining ethnic conflict in a neighbourhood. Blair, (2023) found that whether white agents departed from a neighbourhood was related to the cost of outside options, including that higher desire to be near other white agents could prevent an agent from leaving the neighbourhood as it integrates, rather than departing as is commonly expected.
@@
We are interested in expanding Schelling segregation to a social network modelled with an RDPG. Social networks are less dependent on geographic proximity than they have been in the past, since people are able to form connections online rather than only in person, so a person’s social network may not match the level of integration in e.g. their residential neighbourhood. As such, we give particular attention to the following two papers which use a network graph in this way. Henry et al., (2011) use a Markov chain to generate a stochastic sequence of graphs comprised of actors (nodes) connected by links (edges). The model evolves based on Schelling-based assessments of the attribute distance between two actors connected by a given edge, with links being terminated and rewired to prioritise connections between actors with smaller attribute distances. They use networks of 100 actors connected by 300 edges, and find that segregation emerges in the graphs in similar ways to spatial Schelling models. Similarly, Gretha et al., (2018) create graphs of 1000 agents with 4 initial links, and base network evolution behaviours on a tolerance of 50%. They find that segregation emerges in the network that is greater than the amount desired by the agents, similar to in Schelling’s original results (though this may not be surprising given that this effect has been observed with tolerances as low as 37.5%). Curiously, random behaviour in the selection of which agent to make a new link with caused cascades of dissatisfaction and more movement within the network, whereas ”discriminatory” selection of an agent in the same group caused fewer cascades – an important result for understanding how the dynamics of networks are different to those of spatial models. Based on these results we believe that there is scope for new research that uses much larger and more densely connected graphs, in order to model an agent’s social life more extensively. One of our aims with this paper is to address this gap.
@@
Secondly, we systematically vary both the tolerance of the two groups and their proportions. The combination of these variables has been found to be very important to the Schelling model, since it is the driver of Fossett, (2006)’s ”paradox of weak minority preferences”. While a social network is not subject to the spatial considerations that have previously applied to these results, it is still possible that this effect will occur a social network model, as minority group members may become inundated with links to majority group members, making it difficult to achieve the desired ratio of same- and different-group connections. We will test this by setting the group sizes to 15% and 85% and making small changes to $t$ for both groups.
Finally, we measure $\hat{d}$ alongside the ratios of same- and different-group connections for each agent, as a further investigation of whether it is a good measurement of polarisation in social networks. Our preliminary results using Stochastic Block Models indicate that $\hat{d}$ increases when there are more connections between two groups – i.e. the groups are less polarised – and that it decreases sharply when one group is much larger than the other. Since existing research on the Schelling model has found that segregation and agent unhappiness are both high when one group is much larger than the other, we are interested to see what results we get by measuring $\hat{d}$ for the networks generated in our experiment.
@@ 1.1 Agent-based model
We designed an agent-based model in Julia Datseris et al., (2022) Bezanson et al., (2017) that uses the preference rules of Schelling segregation to change the networking behaviour of the agents. It takes parameters $t\in[0,1]$ for tolerance ($t_{1}\in[0,1]$ and $t_{2}\in[0,1]$ when the tolerance is different between the groups), i.e. the proportion of same-group neighbours needed for the agent to be happy, and $s\in[0,1]$ for the size of the smaller group. The model is initialised with 1000 agents split between 2 groups, who each make 80 random connections on a simple weighted network graph. At each step, each agent evaluates how many of its network neighbours are in the same group as it. If the proportion of same-group neighbours is equal to or above its tolerance $t$, it is happy and does nothing. If the proportion of same-group neighbours is below $t$, it breaks a connection with a neighbour of the opposite group. If at any point the number of total neighbours is lower than 40, it adds a neighbour at random. The model runs for 1000 steps or until all agents are happy. All models were repeated 100 times for each combination of parameters.
@@ 1.2 Network modelling
In order to measure the embedded dimensionality $\hat{d}$ we model the simple weighted network graphs as Random Dot-Product Graphs (RDPGs) Athreya et al., (2017). This graph embedding is chosen because its optimal embedded dimensionality is established before the analysis and is independent of the network size. For a full discussion of the use of RDPG, see Anastasi and Dalla Riva, 2025a .
@@ 1.3 Embedding dimension
We define the dimension of a social network as the optimal choice of $d$ for the RDPG embedding of the network and denote it $\hat{d}$.
@@
An RDPG where every pair of nodes have an identical and independent probability of establishing a connection will have a $\hat{d}$ of 1, since 1 dimension is required for the nodes to make connections at all. It will not have a higher dimensionality because there is nothing else affecting how the edges are distributed. As such, instead of observing a high $\hat{d}$ which decreases as the network becomes segregated, as we have found in our previous work, the network will have a $\hat{d}$ of 1 until it is segregated enough that this noticeably affects whether edges are able to form between the two groups, at which point $\hat{d}$ will increase to 2. We note this because in our previous work we have used Stochastic Block Models, which have enough structure to have a higher $\hat{d}$, whereas the random simple weighted graphs used in this paper are much more random and therefore start with a $\hat{d}$ of 1.
@@ 2.3 Changing group size
To investigate the effect of changing group size, we ran the same model with symmetrical tolerances of $t=0.25$, $t=0.5$, $t=0.85$, and $t=0.95$ so that group size effects were captured when the model was polarised and when it was not. For each tolerance, a minority group size $0.05\leq s\leq 0.5$ was tested in increments of 0.05. We could not test $s\leq 0.05$ due to the minimum of 40 neighbours set for the agent; since $s=0.01$ would give only 10 agents in the minority, they would be structurally incapable of being happy if they have $t_{1}\geq 0.25$ and the model would be unable to stabilise at all. We measured the number of steps it took for all agents to become happy (fig6), the proportion of similar neighbours to total neighbours (fig7), and $\hat{d}$, with full results reported in Table 3.